OpenStudy (anonymous):

Two medians AD and BE of triangle ABC intersect at G at right angles . If AD = 9 cm , BE = 12 cm , find length of AB?

OpenStudy (jmark):

Medians of a triangle intersect in the ratio of 2:1 Length of AB = √(BD)2+(AD)2...... (1) To find BD: AD =AG:GD = 6:3 BE = BG:GE = 8:4 from triangle BDG BD = √BG2+DG2.....(2) = √82+32 =√64+9 = √73 assign value in equation (1) AB = √(√73)2+(9)2 = √73+81 =√154 ICSE class 4 maths question paper - http://icse.edurite.com/icse-question-papers/icse-class-4-maths-question-paper.html